How many different ID cards can be made if there are 9 digits on a card and no digit can be used more than once? What if digits can be repeated?


1
Expert's answer
2022-03-19T02:39:53-0400

Without repetition: "N=10*9*8*7*6*5*4*3*2=3626800" .

With repetition: "N=10^9=1000000000."